How long will it take a given sum of money (say in rupees) to increase 4 times its present value when compounded half year at 7% rate of interest?


1
Expert's answer
2021-03-17T18:32:55-0400

A = the value of the accrued investment.

P = the principal amount.

R = the annual interest rate.

N = the number of times that interest is compounded per unit t.

= the time the money is invested or borrowed for

A= 4, P=1, R= 0.07 N = 2

T =?

Find "\frac{R}{N} = \frac{0.07}{2} = 0.035"


Formula: "A = P [1+\frac{R}{N} ]^{N\times T}"


"4 = 1 [1+\frac{0.07}{2} ]^{2\times T}"


"[1+0.035]^{2\times T} =4"


"(1.071225) ^{T} =4"

Take logarithms on both sides:

Tlog1.071225 = log4


"T = \frac{log4}{log1.071225}"


"T = \frac{0.6020599913}{0.0298806996} = 20.15\; years" = 20 years and 2 months


It will take approximately 20 years and 2 months for a given sum of money (say in rupees) to increase 4 times its present value